OpenStudy (anonymous):

A spring with a spring constant of 100 pounds per foot is loaded with 1-pound weight and brought to equilibrium. It is then stretched an additional 1 inch and released. Find the equation of motion, the amplitude, and the period. Neglect friction.

OpenStudy (zpupster):

k = 100 lb/ft, m = 1 lb / (32.2 ft/s) = 0.03106 slugs F = -kx mx" = -kx x" + (k/m)x = 0 characteristic equation: r^2 + k/m = 0 r = i*sqrt(k/m) x = Asin(sqrt(k/m)t) + Bcos(sqrt(k/m)t) ω = sqrt(k/m) 2π/T = sqrt(k/m) T = 2π*sqrt(m/k) T = 2π*sqrt(0.03106 slugs / 100 lb/ft) T = 0.1107 s (period) x(0) = 1/12 ft = 0.08333 ft x'(0) = 0 1/12 = Asin(0) + Bcos(0) B = 1/12 = 0.08333 ft x' = Aω*cos(ωt) - Bω*sin(ωt) 0 = Aω*cos(0) - (1/12)ω*sin(0) 0 = Aω A = 0 So B is the amplitude. Equation of motion: x = 0.08333*cos[(2π/0.1107)t]

OpenStudy (anonymous):

I've tried entering that in my online homework...and it says its wrong

OpenStudy (zpupster):

Simple Harmonic Motion: k = 100 lb/ft m = 1 lb/32.2 fps² y = A sin ωt = A sin √(k/m) t A = 1 in ω = √(k/m) = √(100/(1/32.2))= 56.745 rad/s f = ω/(2π) = 56.745/(2π) = 9.031 Hz T = period = 1/f = 1/9.031 = 0.1107 s y(t) = A sin ωt = sin (9.031)t